Circular No. 8055

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003O IN UGC 2798 Doug Rich, Hampden, ME, reports his discovery of an apparent su-

pernova (mag ∼ 17.0) on an unfiltered CCD frame taken with a 0.26-m reflector on Jan. 18.96 UT, with a confirming image (showing the new object at mag 16.9) obtained on Jan. 20.15 by M. Schwartz. SN 2003O is located at α = 3^{h}38^{m}09^{s}.13, δ = +40^{o}58′38′′.0 (equinox 2000.0), which is ∼ 9′′.8 west and 21′′.0 south of the center of UGC 2798. The new object is not visible on the red/blue Palomar Sky Survey digital images or on CCD frames taken by Rich on 2002 Sept. 1.20 and Oct. 15.28 (limiting mag 18.0).

FLARING OBJECT IN ORION CLUSTER G. C. Bower, R. Plambeck, and A. Bolatto, University of California at

Berkeley, report the discovery on Jan. 20 at 86 GHz of a flaring object in the Orion cluster, detected with the Berkeley-Illinois-Maryland Association millimeter-wavelength array at an average flux density of 104 mJy (with evidence that it increased from 40 mJy at Jan. 20.188 UT to 143 mJy at Jan. 20.333). The source was not detected in observations made on 2002 Dec. 23, 2003 Jan. 11, and 14, with 3-σ upper limits of 6, 12, and 15 mJy, respectively. Subarcsecond-resolution data yield a position of α = 5^{h}35^{m}11^{s}.80, δ = −5^{o}21′49′′.2 (equinox 2000.0; error < 0′′.1). The source is positionally coincident with radio source A of Garay et al. (1987, Ap.J. 314, 535) and with infrared source 573 of Hillenbrand and Carpenter (2000, Ap.J. 540, 236). The object may be a deeply embedded massive protostar.

COMET C/2002 Y1 (JUELS-HOLVORCEM) Ephemeris extension to IAUC 8039 (from revised orbital elements on

MPEC 2003-A85; magnitudes from H = 9.0, n = 3):

2003 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Jan. 21 14^{h}25.^{m}27 +31^{o}18.^{′}3 1.215 1.658 97.^{o}3 36.^{o}1 11.1
26 14 46.31 +36 14.3 1.112 1.587 98.2 37.9 10.7
31 15 13.33 +41 52.4 1.024 1.515 97.9 40.1 10.4
Feb. 5 15 49.33 +47 57.2 0.956 1.443 96.0 42.8 10.1
10 16 38.43 +53 51.9 0.912 1.371 92.4 46.0  9.8
15 17 44.13 +58 33.3 0.893 1.300 87.3 49.4  9.6
20 19 03.30 +60 50.4 0.900 1.228 81.0 52.7  9.4
25 20 21.96 +60 15.5 0.929 1.158 74.1 55.3  9.3
Mar. 2 21 26.51 +57 30.1 0.978 1.089 67.1 57.0  9.2
2003 January 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green